# Synthetic fractional flux quanta in a ring of superconducting qubits

### v1

@@ Abstract
A ring of capacitively-coupled transmons threaded by a synthetic magnetic field is studied as a realization of a strongly interacting bosonic system. The synthetic flux is imparted through a specific Floquet modulation scheme based on a suitable periodic sequence of Lorentzian pulses that are known as ‘Levitons’. Such scheme has the advantage to preserve the translation invariance of the system and to work at the qubits sweet spots. We employ this system to demonstrate the concept of fractional values of flux quanta. Although such fractionalization phenomenon was originally predicted for bright solitons in cold atoms, it may be in fact challenging to access with that platform. Here, we show how fractional flux quanta can be read-out in the absorption spectrum of a suitable ’scattering experiment’ in which the qubit ring is driven by microwaves.
@@
Introduction.— When matter is constrained to restricted spatial scales and suitably low temperatures, fundamental and unexpected features dominated by the laws of quantum mechanics may emerge; at the same time, new devices utilizing quantum effects can be conceived. This remarkable virtuous cycle provides a foundational principle of mesoscopic physics, and inspired the definition of quantum technology [1]. The emerging field concerned with ultracold atoms moving in magneto-optical circuits of micrometer spatial scales provides a very good example of a mesoscopic quantum system of artificial matter [2, 3, 4]. In this context, genuine notions of mesoscopic superconductivity like the Josephson effect or persistent currents have been providing inspirations to define new schemes in cold atoms quantum technology. In this paper, we go in the opposite direction, thus closing the ’circle’ of cold atoms and mesoscopic superconductivity: we demonstrate how bright solitons originally formulated for attractive bosonic cold atoms can open new avenues in mesoscopic physics and superconducting circuits technology [5, 6, 7, 8, 9]. Bright solitons in a one dimensional lattice have peculiar properties of stability [10], that are predicted to have a strong impact in interferometry utilizing non-classical states of matter [11]. Here, we focus on a striking property of lattice bright solitons confined in a ring-shaped potential pierced by a synthetic magnetic field: as a result of specific solitonic many-body correlations, the system’s persistent current is characterized by a periodicity reflecting a fractional value with respect to the flux quantum $\Phi_{0}$ obtained in the non-interacting case. Crucially for the logic we adopt in the present work, we note that the fractionalization phenomenon manifests itself in a fractional periodicity of the energies of the system as function of the magnetic field, which, following Leggett, are the analogous of Bloch bands for the problem [12]. In the case of bright solitons, such fractionalization is predicted to scale as the inverse of the number of particles $N_{p}$ in the soliton [13]. Such $N_{p}$ dependence makes it difficult to directly observe the flux fractionalization in cold atoms experiments, since $N_{p}$ is typically large.
@@
Figure 1: a) Schematics of a circuit constituted by a chain of transmons, that realizes an attractive Bose-Hubbard model. b) Floquet modulation protocol, consisting in local transmon frequency modulation.
@@
In this work, we propose a system of coupled superconducting qubits to experimentally observe signatures of the fractionalization phenomenon. Specifically, we consider a one dimensional chain of capacitively coupled transmon qubits [14] arranged in a ring geometry, as depicted in Fig. 1a). The low-energy modes in these systems are plasmons that can be excited by microwave photons and propagate through the capacitors. Such excitations behaves as bosons located at the qubit’s sites, and because of the non-linear inductance of the Josephson junctions in the qubits, they can interact with each other. In contrast to the Cooper pair dynamics occurring in Josephson junctions arrays [15], here the boson-boson interaction are effectively attractive [16, 17]. These systems have been realized experimentally and have aroused considerable interest in the fields of driven-dissipative dynamics [18], many-body localization [19, 20, 21, 22], disordered quantum phases of artificial matter  [23], quantum walks [24, 25, 26], and perfect quantum state transfer [27, 28]. The important observation, here, is that the natural regime in which such systems work is the one of moderate number of bosons $N_{p}$, thus opening the door for the verification of predictions for the attractive Bose-Hubbard dynamics [10, 29, 30, 31] in regimes that might be challenging to reach in other physical platforms.
The other important aspect we would like to point out relates to the synthetic gauge field in a spatially closed qubit network. Because of the capacitive qubit-qubit coupling, such a field cannot induce an actual matter-wave motion as it occur for example in cold atoms systems. Here, we demonstrate that the effective synthetic field can be achieved by means of a suitable Floquet driving [32, 33, 34, 35] of the transmon frequency.
@@
System and methods.— Transmons are weakly anharmonic quantum oscillators that are well described by attractive bosons [36, 17]. The latter consist of collective oscillations in the circuit and the attraction results from the weak anharmonicity provided by the Josephson potential in the regime $E_{J}\gg E_{C}$, with $E_{J}$ the Josephson energy and $E_{C}$ the charging energy. In a ring configuration with $N_{s}$ capacitively-coupled transmons, as schematized in Fig. 1, the bosonic excitations can hop from one transmon to the other and realize a Bose-Hubbard model of attractive bosons in a lattice [23, 37, 38]:
@@
where $a_{N_{s}+1}\equiv a_{1}$, $n_{j}=a^{\dagger}_{j}a_{j}$, $\omega\simeq\sqrt{8E_{C}E_{J}}-E_{C}$, $U\simeq E_{C}$, $J_{0}\simeq(C_{0}/(C_{J}+C_{g}))\sqrt{E_{C}E_{J}/8}$, and we set $\hbar=1$ henceforth. The charging energy of the transmons $E_{C}=e^{2}/2(C_{J}+C_{g})$ is expressed in terms of the capacitance of the Josephson junction $C_{J}$ and the capacitance to ground of the superconducting islands $C_{g}$, with $C_{0}\ll C_{J},C_{g}$ the capacitance between nearest neighbor transmons. Assuming $(C_{0}/(C_{J}+C_{g}))\sqrt{E_{J}/(8E_{C})}\ll 1$, we have $J_{0}\ll U\ll\omega$ (see  [^1] Supplementary Materials for details on the circuit, on the driving protocol, and on the driven-dissipative dynamics. for details on the circuit Hamiltonian), implying strongly attractive interaction. We note that the transmon frequency can be modulated $\omega\to\omega(t)$ by employing the so-called split-junctions, composed by a parallel of Josephson junctions threaded by a time-dependent external flux $\varphi_{x}(t)$. Special working points are the so-called ’sweet spots’, that correspond to maxima or minima of the energy versus $\varphi_{x}$. In this regimes the qubit properties are particularly stable against flux noise [40, 41, 42, 43].
@@
[…]
In the next section, we propose a specific Floquet protocol that leads to generic Peierls phases in $\{0,2\pi\}$ and that can be also employed with the transmons at the sweet spot. Our Floquet scheme employs the so called ‘Leviton’ dynamics.
Synthetic gauge field by Leviton protocol.— Levitons consist of suitable Lorentzian pulses that are characterized by an exponential Fourier spectrum and yield a quantized phase advanced of $2\pi$. Based on such a feature, Levitons have been originally introduced in quantum physics by Levitov as noise-free electronic excitations for edge states in the Quantum Hall effect regimes [54, 55, 56, 57, 58]. Here, we propose Levitons to implement a synthetic gauge field by Floquet dynamics. The protocol consists in modulating a suitable subset of adjacent $N_{m}<N_{s}$ transmons through a sequence of Leviton pulses of period $T$ and width $\tau$
@@
[…]
that can be adjusted by varying the time delay $\delta t=(t_{j+1}-t_{j})$ between pulses in adjacent transmons. This feature can be explicitly seen for $e^{-4\pi\tau/T}\ll 1$ for which we have $Je^{i\gamma}\simeq J_{0}e^{2\pi i\delta t/T}$ (see also Supplemental material). Indeed, Fig. 2a) shows an approximately linearly increasing phase $\gamma$ between two adjacent transmons versus $\delta t$, for different values of the normalized Lorentzian width $\tau/T$ such that $e^{-4\pi\tau/T}\ll 1$. At the same time, the hopping rate $J/J_{0}$ is also shown to be weakly dependent on $\delta t$, as shown in Fig. 2b).
We comment that, due to the linear dependence of the phase $\gamma$ on $\delta t$, modulating all the transmons with time-shifted pulses does not lead to any synthetic flux (since all phases gained cancel out around the ring). In turn, by modulating only a subset of $N_{m}$ transmons a net synthetic flux can be achieved. This choice introduces two “external” links between the modulated and non-modulated transmons, and the associated hopping rate $J^{\prime}$ is suppressed
[…]
@@
Several comments are now in order. i) The Floquet protocol based on Leviton dynamics can interestingly be employed at the transmon sweet spots. To see this, we notice that the minimum of the modulation $\delta\omega(t)$ is $\omega_{c}=(2\pi/T)\tanh(\pi\tau/T)$ and when the transmons are modulated at the sweet spot they naturally generate a ’dc component’ $\omega_{c}$ [39]. ii) In order to utilize the series of Levitons at the sweet spots, it is then necessary to shift the frequency $\omega$ of the modulated transmons by $\omega_{c}$, so to compensates the dc component. The proper value of the Floquet period $T$ and the width of the Lorentzians $\tau$ can then be adjusted a posteriori, so to match the step $\omega_{c}$ and the bare hopping $J_{0}^{\prime}$. iii) We note that the rescaling of $J_{0}^{\prime}$ breaks the original translation invariance of the Hamiltonian $H_{0}$ and can introduce coupling to higher Floquet bands in the exact time-dependent dynamics. Nevertheless, for sufficiently high Floquet frequency such that $2\pi/T\gg J_{0}^{\prime}e^{2\pi\tau/T}$ these transitions can be suppressed [39]. Finally iv) we point out that requiring the total amplitude of the modulation to be a fraction of the transmon frequency sets the correct hierarchy of frequency scales: $2\pi/T<2\pi/\tau<\omega$. This is not in contrast with the requirement that the Floquet frequency is effectively the higher scale in the system as in each sector of constant number of bosons the dynamics is controlled by $J/U$; typical values of $J/\omega\sim 10^{-2}$ can fully justify a Floquet regime in which $J\ll 2\pi/T\ll\omega$. The numerical comparison between the target and the actual dynamics of the system is carried out in the Supplemental material.
@@
Figure 2: a) Phase and b) modulus of the complex hopping between two transmons modulated through Levitons, as a function of the time delay $\delta t/T$ for four different values of the Lorentzian width $\tau/T$ [in a), only the curve for the smallest width can be distinguished from the other three curves].
@@
Spectrum measurement.— Having established a protocol for imparting a synthetic flux to the system, we now propose an experimentally feasible scheme for detecting the effect of $\Phi$ on the superconducting circuit. Specifically, we consider a transmission line capacitively coupled to the ring of transmons. The ring network can then be coherently driven by microwaves of frequency $\omega_{d}$ and by the same transmission line the outgoing photons can be monitored. This way, we can map-out the spectrum of the system through the amplitude of the reflected wave $S(\omega_{d})$. Following  [17] and by time-averaging the combined Floquet modulation and external driving dynamics (in the rotating frame of the driving), the effective Hamiltonian is modified as [39]
$H_{\rm eff}\to H_{\rm eff}+\frac{\Omega}{2}(a_{1}+a_{1}^{\dagger}),$ (7)
@@
together with $\omega\to\omega-\omega_{d}$, and we have neglected fast rotating terms for $\Omega\ll\omega,\omega_{d}$. Clearly, the coherent driving implies non-number-conserving process (it couples sectors with different number of bosons): The system coherently absorbs photons from the transmission line and incoherently loses photon both via the transmission line and by relaxation processes in each transmon; in addition, dephasing in each transmon can also result in a loss of coherence. We expect that solitonic bands with different $N_{p}$ are separated in energy by $\sim U$. Therefore, in order to avoid band mixing, we require $\Omega\ll U$. For simplicity, we assume that only transmons that are not modulated through the Floquet protocol are connected to the external drive.
@@
We study the evolution of the entire system through a Lindblad master equation [59]
$\frac{\partial\hat{\rho}}{\partial t}={\cal L}[\hat{\rho}]=-i[H,\hat{\rho}]+{\cal D}_{1}[\hat{\rho}]+{\cal D}_{\phi}[\hat{\rho}],$ (8)
@@
with ${\cal D}_{\alpha}[\hat{\rho}]=\sum_{j}\left[O_{j,(\alpha)}\hat{\rho}O_{j,(\alpha)}^{\dagger}-\frac{1}{2}\{O^{\dagger}_{j,(\alpha)}O_{j,(\alpha)},\hat{\rho}\}\right]$ a general dissipator, that for relaxation processes is specified by $O_{j,(1)}=\sqrt{\gamma_{1}}a_{j}$ and for dephasing processes is described by $O_{j,(\phi)}=\sqrt{\gamma_{\phi}}a^{\dagger}_{j}a_{j}$, with $\gamma_{1}$ and $\gamma_{\phi}$ the transmon relaxation and pure dephasing rates, respectively. To model the system output, we follow Refs. [36, 17] and assume that the site $1$ couples with rate $\Gamma$ to the transmission line. Following a quantum Langevin approach [60, 61, 62], coherent irradiation results in the input field mode amplitude being related to the driving strength $\Omega$ as $\sqrt{\Gamma}\langle a^{\dagger}_{\rm in}\rangle\approx i\Omega$ and the output field is $a^{\dagger}_{\rm out}\approx\sqrt{\Gamma}a^{\dagger}_{1}$. In this regime, photon reflection amplitude is then expressed as
@@
where $\hat{\rho}_{ss}$ is the steady-state density matrix satisfying $\partial\hat{\rho}_{ss}/\partial t=0$. We then perform numerical simulations for system of size $N_{s}=4$, keeping up to $N_{p}=3$ excitations, and we choose experimentally relevant parameters taken from Ref. [17]: $U=0.188~{}{\rm GHz}$, $J=0.041~{}{\rm GHz}$, $\Gamma=16.0~{}{\rm MHz}$, $\gamma_{1}=0.01~{}{\rm MHz}$.
@@
Figure 3: a) Modulus of the wave reflected off the driven system as a function of the synthetic flux $\Phi$, detuning $\omega_{d}-\omega$ (in GHz), and driving amplitude $\Omega=0.01$ GHz, for a chain of $N_{s}=4$ transmons. A clear single-particle spectrum ($N_{p}=1$) appears as a broad bright signal around zero detuning and shows the expected $2\pi$ periodicity as a function of the synthetic flux. At negative detuning, a band made of two-particle bound states characterizing bright solitons appears with the predicted periodicity halved to $\pi$ (lower part of the spectra). b) For further increasing the external drive $\Omega=0.04$ GHz three-particle bound states emerge with one third periodicity. c) Superimposed spectra of the $N_{p}=1,2,3$ sectors, showing the correspondence with the output signal. d) Zoom of the $N_{p}=3$ sector. The rest of the system parameters are given in the main text. The color bar in a) is made non-linear by setting $z(x,y)^{2/5}$.

### v2

@@ Abstract
A ring of capacitively coupled transmons threaded by a synthetic magnetic field is studied as a realization of a strongly interacting bosonic system. The synthetic flux is imparted through a specific Floquet modulation scheme based on a suitable periodic sequence of Lorentzian pulses that are known as “Levitons.” Such scheme has the advantage to preserve the translation invariance of the system and to work at the qubit sweet spots. We employ this system to demonstrate the concept of fractional values of flux quanta. Although such fractionalization phenomenon was originally predicted for bright solitons in cold atoms, it may be in fact challenging to access with that platform. Here, we show how fractional flux quanta can be read out in the absorption spectrum of a suitable “scattering experiment” in which the qubit ring is driven by microwaves.
I Introduction
Networks of mesoscopic-scale systems at sufficiently low temperature can provide us with artificial quantum matter that can be exploited both to study fundamental aspects of quantum science and to design new quantum devices [1]. To this end, different platforms have been explored, ranging from arrays of quantum dots [2] and trapped ions [3] to superconducting circuits [4] and cold atoms [5, 6]. A fruitful viewpoint adopted so far in the community has been that genuine notions of mesoscopic physics of electronic systems such as the Josephson effect, point contacts, or persistent currents, can provide inspiration to define new concepts in quantum technology [7, 8, 9, 10, 11, 12]. Here, we provide a case study going in the opposite direction: We demonstrate that specific notions originally emerging in current research in quantum technology, as in ultracold atoms systems, can inspire electronic mesoscopic physics. Specifically, we will refer to attractive bosonic cold atoms  [13, 14, 15, 16, 17]. On one hand, such systems are difficult to simulate classically. On the other hand, attracting bosons allow the formation of bright solitons that, in turn, bear a great potential in atomtronics and quantum sensing [10, 11, 12]. In particular, bright solitons in a one-dimensional lattice have peculiar properties of stability [18], which are predicted to have a strong impact in interferometry [19].
@@ I Introduction
Here, we focus on a striking property of lattice bright solitons confined in a ring-shaped potential pierced by a synthetic magnetic field: As a result of specific many-body correlations, the system’s persistent current is characterized by a periodicity reflecting a fractional value with respect to the flux quantum $\Phi_{0}$ obtained in the non-interacting case. Crucially for the logic we adopt in the present work, we note that the fractionalization phenomenon manifests itself in a fractional periodicity of the energies of the system as function of the magnetic field, which, following Leggett, are the analogous of Bloch bands for the problem [20]. In the present case, such fractionalization is predicted to scale as the inverse of the number of particles $N_{p}$ in the system [21]. Such $N_{p}$ dependence makes it difficult to directly observe the feature in cold atoms experiments, since $N_{p}$ is typically large.
@@
Figure 1: (a) Schematics of a circuit constituted by a chain of transmons, which realizes an attractive Bose-Hubbard model. (b) Floquet modulation protocol, consisting in local transmon frequency modulation.
@@
In this work, we will define a superconducting qubit quantum simulator for the dynamics of attracting bosons hosted in a ring lattice pierced by an effective magnetic field. In particular, we shall see that our system displays signatures of the fractionalization phenomenon distinctive for bright solitons. Linear chains of capacitivelly coupled qubits have been realized experimentally and have aroused considerable interest in the fields of driven-dissipative dynamics [22], many-body localization [23, 24, 25, 26], disordered quantum phases of artificial matter  [27], quantum walks [28, 29, 30], and perfect quantum state transfer [31, 32]. In these systems, the low-energy modes are plasmons that can be excited by microwave photons and propagate through the capacitors. Such excitations behave as bosons located at the qubit’s sites, and because of the non-linear inductance of the Josephson junctions in the qubits, they can interact with each other. In contrast to the Cooper pair dynamics occurring in Josephson junctions arrays [33], here the boson-boson interaction is effectively attractive [34, 35]. The important observation, here, is that the natural regime in which such systems work is the one of moderate number of bosons $N_{p}$, thus opening the door for analyzing the attractive Bose-Hubbard dynamics [18, 36, 37, 38] in conditions that might be challenging to reach in other physical platforms.
Our capacitively coupled qubits ring system is depicted in Fig. 1(a). In particular, we note that the effective magnetic field cannot induce any actual matter-wave motion as it occur for example in cold atoms. Nevertheless, we can implement a synthetic magnetic field in our superconducting qubit system through a Floquet driving [39, 40, 41, 42, 43] of the transmon frequencies. The latter is suitably engineered to facilitate the operations to be carried out on the superconducting circuit system and protect the transmons performances from decoherence.
II System and methods
@@
Transmons are weakly anharmonic quantum oscillators that are well described by attractive bosons [34, 35]. The latter consist of collective oscillations in the circuit and the attraction results from the weak anharmonicity provided by the Josephson potential in the regime $E_{J}\gg E_{C}$, with $E_{J}$ the Josephson energy and $E_{C}$ the charging energy. In a ring configuration with $N_{s}$ capacitively-coupled transmons, as schematized in Fig. 1, the bosonic excitations can hop from one transmon to the other and realize a Bose-Hubbard model of attractive bosons in a lattice [27, 44, 45]:
@@
where $a_{N_{s}+1}\equiv a_{1}$, $n_{j}=a^{\dagger}_{j}a_{j}$, $\omega\simeq\sqrt{8E_{C}E_{J}}-E_{C}$, $U\simeq E_{C}$, $J_{0}\simeq(C_{0}/(C_{J}+C_{g}))\sqrt{E_{C}E_{J}/8}$, and we set $\hbar=1$ henceforth. The charging energy of the transmons $E_{C}=e^{2}/2(C_{J}+C_{g})$ is expressed in terms of the capacitance of the Josephson junction $C_{J}$ and the capacitance to ground of the superconducting islands $C_{g}$, with $C_{0}\ll C_{J},C_{g}$ being the capacitance between nearest neighbor transmons. Assuming $(C_{0}/(C_{J}+C_{g}))\sqrt{E_{J}/(8E_{C})}\ll 1$, we have $J_{0}\ll U\ll\omega$ (see  [^1] Supplementary Materials for details on the circuit, on the driving protocol, and on the driven-dissipative dynamics. for details on the circuit Hamiltonian), implying strongly attractive interaction. We note that the transmon frequency can be modulated $\omega\to\omega(t)$ by employing the so-called split-junctions, composed by a parallel of Josephson junctions threaded by a time-dependent external flux $\varphi_{x}(t)$. Special working points are the so-called ’sweet spots’, that correspond to maxima or minima of the energy versus $\varphi_{x}$. In this regimes the qubit properties are particularly stable against flux noise [47, 48, 4, 49].
@@
We can perform the unitary transformation $U(t)=\exp\left[i\sum_{j}n_{j}\int^{t}dt^{\prime}\delta\omega_{j}(t^{\prime})\right]$ so that the effective Floquet Hamiltonian in the fundamental band is given by the time-averaged Hamiltonian $H_{\rm eff}=\left\langle U^{\dagger}(t)H(t)U(t)-iU^{\dagger}(t)\dot{U}(t)\right\rangle_{T}$, with $\langle\ldots\rangle_{T}\equiv\int_{0}^{T}\frac{dt}{T}\ldots$. The goal of the protocol is that $H_{\rm eff}$ acquires a complex hopping amplitude ${J_{0}\to J_{0}e^{i\frac{\Phi}{N_{s}}}}$, accounting for the Peierls substitution employed in one-dimensional matterwave circuits in presence of a gauge field [50]. A finite Peierls phase has been achieved in two dimensional arrays of transmons in the hard-core boson limit ($U\to\infty$) with a combination of static gradients and local sinusoidal modulations of $\omega$ that match the gradient steps [23, 51, 52]. We notice that such protocol typically lifts the transmons away from their sweet spot and that in the absence of static gradients, a sinusoidal driving has been demonstrated to yield generic values of Peierls phases in $\{0,2\pi\}$ at the price of suitably coloring the modulation with more than one frequency [53, 54, 55, 39, 56, 57, 58, 59, 60]. We point out that synthetic gauge fields have been also realized in systems of transmons by suitable sets of two-qubit gates in the hard-core boson limit [61, 62]. However, such investigations do not access the bright soliton physics determined by interactions.
In the next section, we propose a specific Floquet protocol that leads to generic Peierls phases in $\{0,2\pi\}$ and that can be also employed with the transmons at the sweet spot. Our Floquet scheme employs the so called Leviton dynamics.
III Synthetic gauge field by Leviton protocol
Levitons consist of suitable Lorentzian pulses that are characterized by an exponential Fourier spectrum and yield a quantized phase advanced of $2\pi$. Based on such a feature, Levitons have been originally introduced in quantum physics by Levitov as noise-free electronic excitations for edge states in the Quantum Hall effect regimes [63, 64, 65, 66, 67]. Here, we propose Levitons to implement a synthetic gauge field by Floquet dynamics. The protocol consists in modulating a suitable subset of adjacent $N_{m}<N_{s}$ transmons through a sequence of Leviton pulses of period $T$ and width $\tau$
@@ III Synthetic gauge field by Leviton protocol
$Je^{-i\gamma}=\langle J_{0}e^{-i\int^{t}dt^{\prime}(\delta\omega(t^{\prime})-\delta\omega(t^{\prime}-\delta t))}\rangle_{T},$ (4)
that can be adjusted by varying the time delay $\delta t=(t_{j+1}-t_{j})$ between pulses in adjacent transmons. This feature can be explicitly seen for $e^{-4\pi\tau/T}\ll 1$ for which we have $Je^{-i\gamma}\simeq J_{0}e^{-2\pi i\delta t/T}$ (see also Supplemental material). Indeed, Fig. 2(a) shows an approximately linearly increasing phase $\gamma$ between two adjacent transmons versus $\delta t$, for different values of the normalized Lorentzian width $\tau/T$ such that $e^{-4\pi\tau/T}\ll 1$. At the same time, the hopping rate $J/J_{0}$ is also shown to be weakly dependent on $\delta t$, as shown in Fig. 2(b).
We comment that, due to the linear dependence of the phase $\gamma$ on $\delta t$, modulating all the transmons with time-shifted pulses does not lead to any synthetic flux (since all phases gained cancel out around the ring). In turn, by modulating only a subset of $N_{m}$ transmons a net synthetic flux can be achieved. This choice introduces two “external” links between the modulated and non-modulated transmons, and the associated hopping rate $J^{\prime}$ is suppressed:
$J^{\prime}=\langle J_{0}e^{-i\int^{t}dt^{\prime}\delta\omega(t^{\prime}-t_{j})}\rangle_{T}=-J_{0}e^{-2\pi\tau/T}.$ (5)
@@
Several comments are now in order. (1) The Floquet protocol based on Leviton dynamics can interestingly be employed at the transmon sweet spots. To see this, we notice that the minimum of the modulation $\delta\omega(t)$ is $\omega_{c}=(2\pi/T)\tanh(\pi\tau/T)$ and when the transmons are modulated at the sweet spot they naturally generate a ’dc component’ $\omega_{c}$ [46]. (2) In order to utilize the series of Levitons at the sweet spots, it is then necessary to shift the frequency $\omega$ of the modulated transmons by $\omega_{c}$, so to compensates the dc component. The proper value of the Floquet period $T$ and the width of the Lorentzians $\tau$ can then be adjusted a posteriori, so to match the step $\omega_{c}$ and the bare hopping $J_{0}^{\prime}$. (3) We note that the rescaling of $J_{0}^{\prime}$ breaks the original translation invariance of the Hamiltonian $H_{0}$ and can introduce coupling to higher Floquet bands in the exact time-dependent dynamics. Nevertheless, for sufficiently high Floquet frequency such that $2\pi/T\gg J_{0}^{\prime}e^{2\pi\tau/T}$ these transitions can be suppressed [46]. (4) Finally, we point out that requiring the total amplitude of the modulation to be a fraction of the transmon frequency sets the correct hierarchy of frequency scales: $2\pi/T<2\pi/\tau<\omega$. This is not in contrast with the requirement that the Floquet frequency is effectively the higher scale in the system as in each sector of constant number of bosons the dynamics is controlled by $J/U$; typical values of $J/\omega\sim 10^{-2}$ can fully justify a Floquet regime in which $J\ll 2\pi/T\ll\omega$. The numerical comparison between the target and the actual dynamics of the system is carried out in the Supplemental material [46].
@@
Figure 2: (a) Phase and (b) modulus of the complex hopping between two transmons modulated through Levitons, as a function of the time delay $\delta t/T$ for four different values of the Lorentzian width $\tau/T$ [in (a), only the curve for the smallest width can be distinguished from the other three curves].
IV Spectrum measurement
@@
Having established a protocol for imparting a synthetic flux to the system, we now propose an experimentally feasible scheme for detecting the effect of $\Phi$ on the superconducting circuit. Specifically, we consider a transmission line capacitively coupled to the ring of transmons and add the driving term $\Omega\cos(\omega_{d}t)(a_{1}+a_{1}^{\dagger})$ to the Hamiltonian, where $\Omega$ characterizes the strength of the driving. The ring network can then be coherently driven by microwaves of frequency $\omega_{d}$ and by the same transmission line the outgoing photons can be monitored. This way, we can map-out the spectrum of the system through the amplitude of the reflected wave $S(\omega_{d})$. Following Ref. [35] and by time averaging the combined Floquet modulation and external driving dynamics (in the rotating frame of the driving), the effective Hamiltonian becomes [46]
$H_{\rm eff}$ $=$ $\sum_{j=1}^{N_{s}}\left[(\omega-\omega_{d})n_{j}-\frac{U}{2}n_{j}(n_{j}-1)\right]$ (7)
- $J_{0}\sum_{j=1}^{N_{s}}(e^{i\Phi/N_{s}}a^{\dagger}_{j+1}a_{j}+{\rm H.c.})+\frac{\Omega}{2}(a_{1}+a_{1}^{\dagger}),$
@@
and we have neglected fast rotating terms for $\Omega\ll\omega,\,\omega_{d}$. Clearly, the coherent driving implies non-number-conserving process (it couples sectors with different number of bosons): The system coherently absorbs photons from the transmission line and incoherently loses photons both via the transmission line and by relaxation processes in each transmon; in addition, dephasing in each transmon can also result in a loss of coherence. We expect that solitonic bands with different $N_{p}$ are separated in energy by $\sim U$. Therefore, in order to avoid band mixing, we require $\Omega\ll U$. For simplicity, we assume that only transmons that are not modulated through the Floquet protocol are connected to the external drive.
@@
We study the evolution of the entire system phenomenologically through a Lindblad master equation [68, 69, 70, 71]
$\frac{\partial\hat{\rho}}{\partial t}={\cal L}[\hat{\rho}]=-i[H_{\rm eff},\hat{\rho}]+{\cal D}_{1}[\hat{\rho}]+{\cal D}_{\phi}[\hat{\rho}$], (8)
@@
with ${\cal D}_{\alpha}[\hat{\rho}]=\sum_{j}\left[O_{j,(\alpha)}\hat{\rho}O_{j,(\alpha)}^{\dagger}-\frac{1}{2}\{O^{\dagger}_{j,(\alpha)}O_{j,(\alpha)},\hat{\rho}\}\right$] a general dissipator, which for relaxation processes is specified by $O_{j,(1)}=\sqrt{\gamma_{1}}a_{j}$ and for dephasing processes is described by $O_{j,(\phi)}=\sqrt{\gamma_{\phi}}a^{\dagger}_{j}a_{j}$, with $\gamma_{1}$ and $\gamma_{\phi}$ the transmon relaxation and pure dephasing rates, respectively, which within a phenomenological approach should be interpreted as effective ones. To model the system output, we follow Refs. [34, 69, 35] and assume that the site $1$ couples with rate $\Gamma$ to the transmission line. Following a quantum Langevin approach [72, 68, 73], coherent irradiation results in the input field mode amplitude being related to the driving strength $\Omega$ as $\sqrt{\Gamma}\langle a^{\dagger}_{\rm in}\rangle\approx i\Omega$ and the output field is $a^{\dagger}_{\rm out}\approx\sqrt{\Gamma}a^{\dagger}_{1}$. In this regime, photon reflection amplitude is then expressed as
@@
where $\hat{\rho}_{ss}$ is the steady-state density matrix satisfying $\partial\hat{\rho}_{ss}/\partial t=0$. We then perform numerical simulations for system of size $N_{s}=4$, keeping up to $N_{p}=3$ excitations, and we choose experimentally relevant parameters taken from Ref. [35]: $\omega=3.9~{\rm GHz}$, $U=0.188~{\rm GHz}$, $J=0.041~{\rm GHz}$.
@@
Figure 3: (a) Modulus of the wave reflected off the driven system as a function of the synthetic flux $\Phi$, detuning $\omega_{d}-\omega$ (in GHz), and driving amplitude $\Omega=0.01$ GHz, for a chain of $N_{s}=4$ transmons. A clear single-particle spectrum ($N_{p}=1$) appears as a broad bright signal around zero detuning and shows the expected $2\pi$ periodicity as a function of the synthetic flux. At negative detuning, a band made of two-particle bound states characterizing bright solitons appears with the predicted periodicity halved to $\pi$ (lower part of the spectra). (b) For further increasing the external drive $\Omega=0.04$ GHz three-particle bound states emerge with one third periodicity. (c) Superimposed spectra of the $N_{p}=1,2,3$ sectors, showing the correspondence with the output signal. (d) Zoom of the $N_{p}=3$ sector. The rest of the system parameters are given in the main text. The color bar in a) is made non-linear by setting $z(x,y)^{2/5}$.